# From Mono- to Hexa-Interstitials: Computational Insights into Carbon Defects in Diamond

### v1

@@ Abstract
We present a comprehensive first-principles investigation of carbon self-interstitial defects in diamond, ranging from mono- to hexa-interstitial complexes. By quantum mechanical density functional theory, empowered by interatomic potential models, we efficiently sample the complex configurational landscape and identify both known and previously unreported defect geometries. Our results reveal a pronounced energetic driving force for aggregation: the formation energy per interstitial decreases systematically from isolated split interstitials to compact multi-interstitial clusters, with the tetra-interstitial platelet emerging as a particularly stable structural motif. Additionally, charge analysis indicates that the predominantly covalent bonding in diamond becomes more polar within the defect centers. Analysis of defect energy levels shows that only the investigated mono-, di-, penta-, and hexa-interstitial complexes introduce in-gap electronic states, whereas the tri- and tetra-interstitial clusters are electronically inert. Vibrational spectroscopies further reveal that self-interstitials generate characteristic signatures. Short carbon–carbon bonds inside the defect cores give rise to high-frequency vibrational modes between 1375 and 1925 cm-1, which are strongly IR-active but exhibit weak Raman activity. Taken together, these findings provide a coherent picture of the structural, electronic, and vibrational characteristics of carbon self-interstitials and establish a robust framework for their experimental identification.
@@ I Introduction
In this work, we perform a systematic first-principles investigation of carbon self-interstitial defects. Motivated by suggestions that TR12 may involve aggregates of up to six carbon self-interstitials [35], we extend our analysis to clusters containing one to six self-interstitial atoms. We first investigate the structural properties of these configurations and then compare their relative formation energies to identify the most stable defects. We also investigate the electronic and vibrational properties of the most stable complexes, providing Raman and IR spectroscopic fingerprints that enable the identification of electronically and optically silent configurations.
@@ II Methodology
Defect model systems were constructed in a $4\times 4\times 4$ cubic supercell containing 512 carbon atoms, derived from the conventional diamond unit cell. With this supercell size, Brillouin-zone sampling at the $\Gamma$ point provides converged defect energetics and electronic properties. During structural relaxations, the lattice vectors were kept fixed to isolate local distortions induced by the defects and to avoid spurious cell-wide strain effects.
To easily generate more realistic initial configurations for carbon self-interstitial clusters in diamond, we employed a geometric Frenkel-pair construction carried out directly in configuration space. A spherical cavity of radius $R$ was created at the center of a 512 atom diamond supercell by removing all atoms within a distance $R$ from the cell center. The cavity radii (1.6–3.9 Å) were selected based on the first several coordination shells of diamond, enabling a systematic modeling of voids capable of accommodating mono- to multi-interstitial complexes. Within this cavity, $m$ additional carbon atoms ($1\leq m\leq 6$) were inserted using rejection Monte-Carlo sampling with a hard-sphere exclusion distance of 1.2 Å, ensuring physically reasonable C-C separations. Candidate positions were drawn from an isotropic, volume-uniform distribution, allowing the initial aggregates to sample a broad range of compact and extended motifs. For each cluster size and $R$, 200 statistically independent initial structures were generated.
@@
To generate continuous spectra, all discrete mode intensities were convoluted using Gaussian functions with a broadening of 5 cm-1.
@@ III.1 Structural Properties
Our HSE06 calculations indicate that the neutral $I_{1}^{\langle 001\rangle}$ in diamond has a singlet ($S=0$) ground state, with a triplet ($S=1$) state lying approximately $40$ meV higher in energy. These findings agree with the previous theoretical studies [31, 30, 29]. In addition, experimental measurements have identified a defect, known as the R2 center, exhibiting EPR activity, which has recently been shown through ZFS calculations to correspond to the $I_{1}^{\langle 001\rangle}$ configuration [33]. Its temperature-dependent EPR intensities place the triplet roughly $50$ meV above the singlet ground state, in close agreement with our result [48].
We find that the structural symmetry of $I_{1}^{\langle 001\rangle}$ exhibits a pronounced dependence on both the xc-functional and the spin multiplicity, as summarized in Table 1. Using semilocal functionals, namely the local-density approximation (LDA) [27] and PBE, we obtain the triplet $D_{2d}$ configuration as the ground state, with the corresponding triplet in $D_{2}$ symmetry lying approximately 0.5 eV higher in energy, in contrast to previous results [31, 30]. Conversely, when the metastable singlet is taken as the reference ground state, a Jahn–Teller distortion is recovered: the singlet in $D_{2}$ symmetry is 0.47 eV lower in energy than its $D_{2d}$ counterpart, in good agreement with the value reported by Goss et al. [31].
Table 1: Relative energies (eV) of the $I_{1}^{\langle 001\rangle}$ configuration for different symmetries and spin states obtained with LDA, PBE, and HSE06. The lowest-energy configuration for each functional is taken as the zero reference.
@@
$D_{2}$ singlet 0.02 0.35 0.27
$D_{2}$ triplet 0.48 0.49
$D_{2d}$ singlet 0.49 0.80 0.00
$D_{2d}$ triplet 0.00 0.00 0.04
Within the HSE06 functional, the $D_{2}$ geometry is not even metastable in the triplet state; upon relaxation, it spontaneously transforms to the higher-symmetry $D_{2d}$ configuration, indicating that $D_{2d}$ is the only stable minimum at this spin multiplicity. By contrast, the singlet state is stable in both symmetries, but lies 0.27 eV lower in energy in the $D_{2d}$ geometry than in $D_{2}$.
@@
[…]
@@
For the penta-interstitial, we identify two novel configurations. Configuration $\mathrm{C}^{a}_{\mathrm{5i}}$ consists of five aggregated split-interstitial units, has crystal symmetry $C_{2}$, and a triplet ground state. Configuration $\mathrm{C}^{b}_{\mathrm{5i}}$ lies 3.91 eV higher in energy than $\mathrm{C}^{a}_{\mathrm{5i}}$; it is oriented perpendicular to the (111) plane and is formed by a zigzag stacking of five carbon interstitials occupying hexagonal and tetrahedral interstitial sites. This defect has a $C_{1h}$ symmetry.
Finally, we identify two novel hexa-interstitial configurations, both formed through the aggregation of split-interstitial units. Configuration $\mathrm{C}^{a}_{\mathrm{6i}}$ induces a pronounced lattice distortion and exhibits $C_{1}$ symmetry. This configuration is very similar to structure $\mathrm{C}^{a}_{\mathrm{5i}}$; however, the presence of an additional split-interstitial, arranged so as to break one of the symmetry planes, lowers the symmetry. Configuration $\mathrm{C}^{b}_{\mathrm{6i}}$, which at first glance resembles three parallel di-interstitials of type $\mathrm{C}^{b}_{\mathrm{2i}}$, in fact consists of a combination of one $\mathrm{C}^{b}_{\mathrm{2i}}$ di-interstitial and one $\mathrm{C}^{a}_{\mathrm{4i}}$ tetra-interstitial unit. Two dangling bonds are present in the $\mathrm{C}^{b}_{\mathrm{2i}}$-like Humble segment of this complex. This configuration lies 3.18 eV higher in energy than the most stable hexa-interstitial structure.
@@ III.2 Bond Analysis
Bader charge analysis [55] shows that, across all configurations from $\mathrm{C}^{a}_{\mathrm{1i}}$ to $\mathrm{C}^{a}_{\mathrm{6i}}$, a charge transfer of roughly 1.5 e occurs among the defect carbon atoms, with similar redistributions observed in the remaining systems. This systematic charge imbalance introduces a polar component into the core C-C bonds, partially shifting their character away from purely covalent. The resulting polarization enhances the local electrostatic attraction between the involved atoms, thereby strengthening these bonds relative to the surrounding, more covalent lattice bonds. Consequently, the bonds experiencing the largest charge transfer consistently correspond to the most negative ICOHP values within each defect structure, reflecting their increased bond strength.
@@ III.3 Formation Energies
For the di-interstitial complexes, configuration $\mathrm{C}^{a}_{\mathrm{2i}}$ is stable only in neutral and positive charge states, indicating that it has no empty defect level in the band gap that could accommodate an additional electron. In the neutral state, configurations $\mathrm{C}^{a}_{\mathrm{2i}}$, $\mathrm{C}^{b}_{\mathrm{2i}}$, and $\mathrm{C}^{c}_{\mathrm{2i}}$ have formation energies of 8.2, 8.6, and 8.9 eV per interstitial, respectively. By contrast, configurations $\mathrm{C}^{b}_{\mathrm{2i}}$ and $\mathrm{C}^{c}_{\mathrm{2i}}$ are stable in positive, neutral, and negative charge states. At low and high Fermi-level energies, the positively and negatively charged states of the $\mathrm{C}^{b}_{\mathrm{2i}}$ defect, respectively, are the most stable among the di-interstitial configurations considered, consistent with previous findings [31].
@@
III.4 Electronic Properties
Figure 4: (a) Electronic structure for the ground states of the neutral defects using HSE06 functional. Kohn-Sham levels are represented by spin-up ($\uparrow$) and spin-down ($\downarrow$). The valence band (VB) and conduction band (CB) are depicted in brown and cream, respectively. (b) Representation of Kohn-Sham orbitals of the relevant states. The light cyan and yellow lobes exhibit negative and positive isovalues. The isosurface absolute value is set to $7\times 10^{-7}$ Å-3.

### v2

@@ Abstract
We present a comprehensive first-principles investigation of carbon self-interstitial defects in diamond, ranging from mono- to hexa-interstitial complexes. By quantum mechanical density functional theory, empowered by interatomic potential models, we efficiently sample the complex configurational landscape and identify both known and previously unreported defect geometries. Our results reveal a pronounced energetic driving force for aggregation: the formation energy per interstitial decreases systematically from isolated split interstitials to compact multi-interstitial clusters, with the tetra-interstitial platelet emerging as a particularly stable structural motif. Additionally, charge analysis indicates that the predominantly covalent bonding in diamond becomes more polar within the defect centers. Analysis of defect energy levels shows that only the investigated mono-, di-, penta-, and hexa-interstitial complexes introduce in-gap electronic states, whereas the tri- and tetra-interstitial clusters are electronically inert. Vibrational spectroscopies further reveal that self-interstitials generate characteristic signatures. Short carbon–carbon bonds inside the defect cores give rise to high-frequency vibrational modes between 1375 and 1925 cm-1, which are strongly IR-active but exhibit weak Raman activity. Through a systematic analysis of metastable configurations, we identify the 3H defect center as a neutral di-interstitial defect. Based on this identification, we further suggest that the TR12 center may arise from a 3H-containing defect like a metastable hexa-interstitial configuration. Taken together, these findings provide a coherent picture of the structural, electronic, and vibrational characteristics of carbon self-interstitials and establish a robust framework for their experimental identification.
@@ I Introduction
In this work, we perform a systematic first-principles investigation of carbon self-interstitial defects. Motivated by suggestions that TR12 may involve aggregates of up to six carbon self-interstitials [35], we extend our analysis to clusters containing one to six self-interstitial atoms that can still be treated accurately with first-principles methods in a computationally feasible model system. We first investigate the structural properties of these configurations and then compare their relative formation energies to identify the most stable defects. We also investigate the electronic and vibrational properties of the most stable complexes, providing Raman and IR spectroscopic fingerprints that enable the identification of electronically and optically silent configurations. Furthermore, we assess candidate carbon-interstitial-related color centers in the visible range, including the TR12 center, by comparing computed photoluminescence spectra with experiment.
@@ II Methodology
Defect model systems were constructed in a $4\times 4\times 4$ cubic supercell containing 512 carbon atoms, derived from the conventional diamond unit cell. With this supercell size, Brillouin-zone sampling at the $\Gamma$ point provides converged defect energetics and electronic properties. During structural relaxations, the lattice vectors were kept fixed to isolate local distortions induced by the defects and to avoid spurious cell-wide strain effects. Within this framework, the carbon number density increases from $3.522$ to $3.556$ g/cm-3 when going from mono- to hexa-interstitial carbon complexes.
To easily generate more realistic initial configurations for carbon self-interstitial clusters in diamond, we employed a geometric Frenkel-pair construction carried out directly in configuration space. A spherical cavity of radius $R$ was created at the center of a 512 atom diamond supercell by removing all atoms within a distance $R$ from the cell center. The cavity radii (1.6–3.9 Å) were selected based on the first several coordination shells of diamond, enabling a systematic modeling of voids capable of accommodating mono- to multi-interstitial complexes. Within this cavity, $m$ additional carbon atoms ($1\leq m\leq 6$) were inserted using rejection Monte-Carlo sampling with a hard-sphere exclusion distance of 1.2 Å, ensuring physically reasonable C-C separations. Candidate positions were drawn from an isotropic, volume-uniform distribution, allowing the initial aggregates to sample a broad range of compact and extended motifs. For each cluster size and $R$, 200 statistically independent initial structures were generated. By deliberately sampling high-energy, distorted, and non-equilibrium configurations, we ensure adequate exploration of the defect configurational space.
@@
To evaluate the Raman and IR activities according to each vibrational mode, we compute Eqs. 2 and 3 by displacing the atoms along both the positive and negative directions defined by the phonon eigenvectors. This procedure requires 6N66N-6 single-point calculations, where $N$ is the number of atoms in the supercell.
Phonon modes were obtained using the finite-difference method as implemented in VASP. The resulting eigenvectors were mass-scaled to yield normalized displacement vectors. The displacement amplitude was set to $0.015$ Å. To generate continuous spectra, all discrete mode intensities were convoluted using Gaussian functions with a broadening of 5 cm-1.
[…]
$S_{\lambda}=\omega_{\lambda}Q_{\lambda}^{2}\bigg/2\hbar,$ (5)
where $\hbar$ is the reduced Planck constant and the configurational coordinate $Q_{\lambda}$ for an optical process is defined as
$Q_{\lambda}=\sum_{\alpha}\sqrt{m_{\alpha}}\langle(\bf{R}_{e,\alpha}-{\bf{R}}_{g,\alpha})|{\bf{u}}_{\alpha,\lambda}\rangle.$ (6)
Here, ${\bf{R}}_{g}$ and ${\bf{R}}_{e}$ are ground and excited state atomic coordinates, while ${\bf{u}}_{\alpha,\lambda}$ indicates the normalized displacement vector corresponding to mode $\lambda$ with frequency $\omega_{\lambda}$ and $m_{\alpha}$ is mass of atom $\alpha$. The total HRF and $Q^{2}$ are defined as $\rm{HRF}=\sum S_{\lambda}$ and $Q^{2}=\sum Q^{2}_{\lambda}$, respectively. They provide measures for the average number of phonons involved in the emission and the difference between the initial and final state geometries. From the HRF, the DWF can be defined as $\text{DWF}=\exp(-\rm{HRF})$.
The spectral function that underlies the computation of the PSB is
$S(\omega)=\sum_{\lambda}S_{\lambda}\delta(\omega-\omega_{\lambda}),$ (7)
where the $\delta$-function is replaced by Gaussian with a broadening of 3 meV. Once the electron-phonon spectral function is computed, the spectral distribution function can be determined as
$A(E_{\mathrm{ZPL}}/\hbar-\omega)=\int dt\exp(S(0)-S(t))\exp(i\omega t-\gamma|t|),$ (8)
where $S(t)$ is the Fourier transform of $S(\omega)$ and $\gamma$ determines the lifetime broadening of individual contributions. The PL intensity is proportional to $\omega^{3}A(\omega)$.
@@ III.1 Structural Properties
Our HSE06 calculations indicate that the neutral $I_{1}^{\langle 001\rangle}$ in diamond has a singlet ($S=0$) ground state, with a triplet ($S=1$) state lying approximately $20$ meV higher in energy. In fact, this energy difference equals 40 meV when the von Barth procedure is employed [51], which avoids spin-state mixing. These findings agree with the previous theoretical studies [31, 30, 29]. In addition, experimental measurements have identified a defect, known as the R2 center, exhibiting EPR activity, which has recently been shown through ZFS calculations to correspond to the $I_{1}^{\langle 001\rangle}$ configuration [33]. Its temperature-dependent EPR intensities place the triplet roughly $50$ meV above the singlet ground state, in close agreement with our result [52].
We find that the structural symmetry of $I_{1}^{\langle 001\rangle}$ exhibits a pronounced dependence on both the xc-functional and the spin multiplicity, as summarized in Table 1. Using semilocal functionals, namely the local-density approximation (LDA) [27] and PBE, we obtain the triplet $D_{2d}$ configuration as the ground state. In contrast, HSE06 stabilizes a singlet state and consequently predicts the triplet configuration to be metastable. When switching from the stable to the metastable configuration, the symmetry changes to $D_{2}$ in PBE and LDA calculations, whereas in HSE06 the structure remains in the $D_{2d}$ symmetry. In the former cases, this behavior can be interpreted as a Jahn–Teller distortion: the singlet state with $D_{2}$ symmetry is about 0.45 eV lower in energy than its $D_{2d}$ counterpart, in good agreement with the value reported by Goss et al. [31]. It is noteworthy that in HSE06 calculations the $D_{2d}$ geometry is 0.27 eV lower in energy than the $D_{2}$ configuration.
Table 1: Relative energies (eV) of the $I_{1}^{\langle 001\rangle}$ configuration for different symmetries and spin states obtained with LDA, PBE, and HSE06. The lowest-energy configuration for each functional is taken as the zero reference. The value reported in parentheses is calculated within the von Barth procedure.
@@
$D_{2}$ singlet 0.02 (0.04) 0.33 (0.67) 0.27(0.55)
$D_{2}$ triplet
$D_{2d}$ singlet 0.49 (0.97) 0.79 (1.59) 0.00
$D_{2d}$ triplet 0.00 0.00 0.02 (0.04)
@@
Three 2Ci configurations, known as $\pi$-bonded, Humble, and di-$\langle 001\rangle$ split-interstitial ($I_{2}^{\langle 001\rangle}$), shown in Fig. 1 as $\mathrm{C}^{a}_{\mathrm{2i}}$-$\mathrm{C}^{c}_{\mathrm{2i}}$, comprise our dataset, listed in order of increasing energy. Further structural details are also provided in Refs. [31, 30, 54]. The $\pi$-bonded di-interstitial, formed by two bond-centered mono-interstitials on opposite sides of the same hexagonal ring, exhibits $C_{2h}$ symmetry and a singlet ground state. The Humble configuration consists of two $I_{1}^{\langle 001\rangle}$ units occupying next-nearest-neighbor sites. This exhibits $C_{2v}$ symmetry and a singlet ground state; the triplet state lies 0.13 eV higher in energy. Notably, the positively charged Humble defect has been proposed to be related to the 3H optical center, with a ZPL energy of 2.46 eV [31]. However, our findings indicate that such a conclusion about the charge state is incorrect because that charge state is not photostable under blue emission (see Sec. III.3). We rather identify 3H optical center with the neutral Humble ($\mathrm{C}^{b}_{\mathrm{2i}}$) defect in Sec. III.6. Finally, the $I_{2}^{\langle 001\rangle}$ consists of two $I_{1}^{\langle 001\rangle}$ defects occupying nearest-neighbor sites, exhibits $C_{2h}$ symmetry, and has historically been attributed to the triplet-state EPR center R1 [54, 33]. In agreement with prior works, we recover the same stability ordering: the $\pi$-bonded configuration is lowest in energy; the Humble one is 0.84 eV higher; and the $I_{2}^{\langle 001\rangle}$ is 1.40 eV higher, all relative to the $\pi$-bonded configuration.
@@
For the penta-interstitial, we identify two novel configurations. Configuration $\mathrm{C}^{a}_{\mathrm{5i}}$ consists of five aggregated split-interstitial units, has crystal symmetry $C_{2}$, and a triplet ground state. The singlet spin-state is nearly degenerate with the triplet state, lying 5 meV lower in energy. Configuration $\mathrm{C}^{b}_{\mathrm{5i}}$ lies 3.91 eV higher in energy than $\mathrm{C}^{a}_{\mathrm{5i}}$; it is oriented perpendicular to the (111) plane and is formed by a zigzag stacking of five carbon interstitials occupying hexagonal and tetrahedral interstitial sites. This defect has a $C_{1h}$ symmetry.
Finally, we identify two novel hexa-interstitial configurations, both formed through the aggregation of split-interstitial units. Configuration $\mathrm{C}^{a}_{\mathrm{6i}}$ induces a pronounced lattice distortion and exhibits $C_{1}$ symmetry. This configuration is very similar to structure $\mathrm{C}^{a}_{\mathrm{5i}}$; however, the presence of an additional split-interstitial, arranged so as to break one of the symmetry planes, lowers the symmetry. Here, the singlet and triplet spin states are nearly degenerate, with the singlet lying within 1 meV below the triplet. Configuration $\mathrm{C}^{b}_{\mathrm{6i}}$, which at first glance resembles three parallel di-interstitials of type $\mathrm{C}^{b}_{\mathrm{2i}}$, in fact consists of a combination of one $\mathrm{C}^{b}_{\mathrm{2i}}$ di-interstitial and one $\mathrm{C}^{a}_{\mathrm{4i}}$ tetra-interstitial unit. Two dangling bonds are present in the $\mathrm{C}^{b}_{\mathrm{2i}}$-like Humble segment of this complex. This configuration lies 3.18 eV higher in energy than the most stable hexa-interstitial structure. The results discussed above are summarized in Table 2. However, singlet–triplet splittings below 10 meV remain inconclusive for identifying the true ground state within the numerical and methodological uncertainty of our approach.
Table 2: The spin multiplicities of defects in the neutral charge state are listed according to the labels used in this work and their conventional names, found in the literature. The triplet-singlet energy difference ($\Delta E_{TS}$) is reported in meV, with negative value indicating a triplet ground state.
Defect labels spin ground state $\Delta E_{ST}$
$\mathrm{C}^{a}_{\mathrm{1i}}$ $\langle 001\rangle$-split singlet 40
$\mathrm{C}^{b}_{\mathrm{1i}}$ Bond-centered triplet 850-850
$\mathrm{C}^{a}_{\mathrm{2i}}$ pipi-bond singlet
$\mathrm{C}^{b}_{\mathrm{2i}}$ Humble singlet 130
$\mathrm{C}^{c}_{\mathrm{2i}}$ di-$\langle 001\rangle$-split triplet 60-60
$\mathrm{C}^{a}_{\mathrm{5i}}$ $\mathrm{C}^{a}_{\mathrm{5i}}$ singlet 5
$\mathrm{C}^{b}_{\mathrm{5i}}$ $\mathrm{C}^{b}_{\mathrm{5i}}$ singlet 140
$\mathrm{C}^{a}_{\mathrm{6i}}$ $\mathrm{C}^{a}_{\mathrm{6i}}$ singlet 1
$\mathrm{C}^{b}_{\mathrm{6i}}$ $\mathrm{C}^{b}_{\mathrm{6i}}$ singlet 60
@@ III.2 Bond Analysis
Bader charge analysis [59] shows that, across all configurations from $\mathrm{C}^{a}_{\mathrm{1i}}$ to $\mathrm{C}^{a}_{\mathrm{6i}}$, a charge transfer of roughly 1.5 electrons occurs among the defect carbon atoms, with similar redistributions observed in the remaining systems. This systematic charge imbalance introduces a polar component into the core C-C bonds, partially shifting their character away from purely covalent. The resulting polarization enhances the local electrostatic attraction between the involved atoms, thereby strengthening these bonds relative to the surrounding, more covalent lattice bonds. Consequently, the bonds experiencing the largest charge transfer consistently correspond to the most negative ICOHP values within each defect structure, reflecting their increased bond strength.
@@ III.3 Formation Energies
For the di-interstitial complexes, configuration $\mathrm{C}^{a}_{\mathrm{2i}}$ is stable only in neutral and positive charge states, indicating that it has no empty defect level in the band gap that could accommodate an additional electron. In the neutral state, configurations $\mathrm{C}^{a}_{\mathrm{2i}}$, $\mathrm{C}^{b}_{\mathrm{2i}}$, and $\mathrm{C}^{c}_{\mathrm{2i}}$ have formation energies of 8.2, 8.6, and 8.9 eV per interstitial, respectively. By contrast, configurations $\mathrm{C}^{b}_{\mathrm{2i}}$ and $\mathrm{C}^{c}_{\mathrm{2i}}$ are stable in positive, neutral, and negative charge states. At low and high Fermi-level energies, the positively and negatively charged states of the $\mathrm{C}^{b}_{\mathrm{2i}}$ defect, respectively, are the most stable among the di-interstitial configurations considered, consistent with previous findings [31]. We note that these defects are photostable under visible illumination only in the neutral charge state: both the positive and negative charge states are photo-converted to neutral via single-photon absorption for photon energies above 1.4 eV. Consequently, the assignment of the positively charged $\mathrm{C}^{b}_{\mathrm{2i}}$ defect to the 3H optical center proposed in Ref. [31], which is excited in the blue, is no longer supported.
@@
III.4 Electronic Structure
Figure 4: (a) Electronic structure for the ground states of the neutral defects using HSE06 functional. Kohn-Sham (KS) energy levels are represented by spin-up ($\uparrow$) and spin-down ($\downarrow$). The valence band (VB) and conduction band (CB) are depicted in brown and cream, respectively. (b) Representation of the KS orbitals of the relevant states. The light cyan and yellow lobes exhibit negative and positive isovalues. The isosurface absolute value is set to $7\times 10^{-7}$ Å-3.